find the equation of the straight line with slope 0.60 that passes through point (5,8).
Answer by Cromlix(4381)

Using y - b =m(x - a)
y - 8 = 0.6(x - 5)
y - 8 = 0.6x - 3
y = 0.6x - 3 + 8
y = 0.6x + 5
Hope this helps.
:-)